Weekly LIPAc highlights (25/Janurary-29/January)

Global Alignment survey

The global alignment survey of the entire LIPAc accelerator from ion source to beam dump cartridge has been completed. The large amount of data recorded is being analysed to assess if further measurements or actions are necessary.

RFQ vacuum system

The RFQ nominal vacuum has been resumed smoothly.

Radio Frequency (RF) System   

The start-up of the RF system with the calibration of the LLRF (Low Level RF) and the RF chains including the debugging started this week. The activity focused on the calibration of the 8 RF chains. Next week the plan is to finalize the calibrations, to check the PSYS interlock after the calibration and check the upgraded anode box. Also, some pending software modifications and LCS updates will be performed (activities for which the RF system ON is needed).

Beam Dump

The installation of the planned upgrade in the deaeration loop (oxygen control loop) of the Beam Dump was completed and water circulation is in preparation.

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.